For each of the following numbers, find the smallest whole number by which it should be multiplied so as to get a perfect square number. Also find the square root of the square number so obtained. 
2028

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student
2028 = 2×2×3×13×13
= (2×2)×(13×13)×3
Here, 3 cannot be paired.
∴ We will multiply 2028 by 3 to get perfect square. New number = 2028×3 = 6084

6084 = 2×2×3×3×13×13

⇒ 6084 = (2×2)×(3×3)×(13×13)

⇒ 6084 =22×32×132

⇒ 6084 = (2×3×13)2

⇒ √6084 = 2×3×13 = 78
Thanks